# Think Outside the Policy: In-Context Steered Policy Optimization

### v2

@@ Abstract
Existing Reinforcement Learning from Verifiable Rewards (RLVR) methods, such as Group Relative Policy Optimization (GRPO), have achieved remarkable progress in improving the reasoning capabilities of Large Reasoning Models (LRMs). However, they exhibit limited exploration due to reliance on on-policy rollouts which are confined to the current policy’s distribution, resulting in narrow trajectory diversity. Recent approaches attempt to expand policy coverage by incorporating trajectories generated from stronger expert models, yet this reliance increases computational cost and such advanced models are often inaccessible. To address these issues, we propose In-Context Steered Policy Optimization (ICPO), a unified framework that leverages the inherent in-context learning capability of LRMs to provide expert guidance using existing datasets. ICPO introduces mixed-policy GRPO with implicit expert forcing, which expands exploration beyond the current policy distribution without requiring advanced LRM trajectories. To further stabilize optimization, ICPO integrates expert region reject sampling to filter unreliable off-policy trajectories and annealed expert-bonus reward shaping to balance early expert guidance with later autonomous improvement. Results demonstrate that ICPO consistently enhances RLVR performance and training stability on mathematical reasoning benchmarks, revealing a scalable and effective RLVR paradigm for LRMs. Our code is available at https://anonymous.4open.science/r/ICPO.
@@
Hsiu-Yuan Huang1,2thanks:   Equal contribution. Work done during internship at Tencent., Chenming Tang1,211footnotemark: 1, Weijie Liu3thanks:   Corresponding author., Clive Bai3, Saiyong Yang3, Yunfang Wu1, [^2] footnotemark: 2 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University 3LLM Department, Tencent ✉  huang.hsiuyuan@stu.pku.edu.cn, wuyf@pku.edu.cn
@@ 3 Preliminary
where $A(\cdot)$ maps demonstrations $\mathcal{D}$ to a task vector $\vartheta=A(\mathcal{D})$ that encodes the expert behavior specific to that task Hendel et al. (2023); Li et al. (2023); Todd et al. (2024); Huang et al. (2024); Liu et al. (2024); Hojel et al. (2024), and $\mathcal{F}(\cdot\,;\vartheta)$ represents the task-specific reasoning function that applies the $\vartheta$ to generate the prediction for query $q$. This leads to the parametric representation:
@@ 5 Experimental Setup
The complete set of hyperparameters used in our training is listed in Appendix B.1. For the main experiments, we use Qwen3-1.7B and Qwen3-8B Yang et al. (2025) as the base models and employ GRPO Shao et al. (2024) as our RL algorithm. We also include Qwen2.5-Math-7B to facilitate comparison with other related methods. We additionally evaluate ICPO on Qwen3-8B-Base and LLaMA-3.1-8B to examine its cross-model generalization. We generate a total of 8 rollout trajectories per prompt. For the on-policy baseline, we use 8 on-policy rollouts. For our mixed-policy GRPO, we follow previous work Yan et al. (2025) and use 1 off-policy rollout and 7 on-policy rollouts to ensure comparability. Evaluation details are provided in Appendix B.2.
@@ 6.1 Main Results
Figure 5: Reward curves of Qwen3-8B over training steps across test (Mean@2) and train sets.
@@ 8 Conclusion
Despite the demonstrated improvements, our approach has several limitations that open avenues for future investigation.
Task Scope.
Our experiments are currently limited to mathematical reasoning tasks under the RLVR framework. Extending ICPO to other domains such as code generation would require additional task-specific adaptation.
Model Capability Assumption.
Our method assumes that the underlying model possesses basic ICL capability; without it, the implicit expert-conditioned rollouts may not yield meaningful supervision. Due to time constraints, we did not conduct experiments on models lacking ICL ability, such as training them to acquire this capability before applying GRPO.
Demonstration Sensitivity.
The model may be sensitive to the choice of ICL demonstrations. We experimented with various heuristic demonstration selection methods, as detailed in the Appendix C. The results show that randomly sampling demonstrations—both across batches and per prompt—yields the best performance. Consequently, we adopt random demonstration selection as the default strategy in our main experiments, without any deliberate curation. We hypothesize that carefully selecting demonstrations tailored to each problem could further improve performance, which we plan to explore in future work.
Expert Data Selection.
Our method eliminates the reliance on stronger external LRMs by leveraging the model’s own ICL capability to treat existing datasets as implicit experts, steering rollouts beyond the default GRPO distribution and enhancing exploration. Although ICPO improves the utility of available high-quality datasets through IEF, it still imposes several constraints on the expert data: (1) expert data must provide complete solutions; (2) the solutions must be of sufficiently high quality to offer effective guidance; and (3) using domain-specific expert data as demonstrations may bias the model toward that domain, potentially degrading its general-purpose capabilities, as detailed in Appendix D.3.
Despite these limitations, ICPO also opens promising directions. First, identifying high-quality, general-purpose expert datasets could further enhance model performance on both in-distribution and out-of-distribution tasks. Second, because IEF encourages the model to adopt the stylistic patterns of expert demonstrations, ICPO may naturally support controllable style transfer in reasoning tasks. We leave these directions to future work.
@@ B.2 Evaluation Settings
We evaluate on six widely used mathematical reasoning benchmarks: AIME24, AIME25, AMC23 Li et al. (2024), Minerva Lewkowycz et al. (2022), Olympiad He et al. (2024), and MATH-500 Hendrycks et al. (2021). For main experiments, the AIME24, AIME25, and AMC23 benchmark, which have relatively small test sets, we report Avg@32 (for Pass@1), while for the other benchmarks we report Pass@1. As a complementary analysis, we also report Pass@32 results in Appendix D.2.
To assess generalization beyond in-domain reasoning, we further test on three out-of-distribution (OOD) benchmarks: ARC-C Clark et al. (2018), GPQA-Diamond Rein et al. (2023), and MMLU-Pro Wang et al. (2024), with multiple-choice options shuffled to prevent contamination. All evaluations are conducted using the LIMO framework Ye et al. (2025). During inference, we follow Yan et al. (2025) and set the generation temperature to 0.6. For ICL evaluation, we randomly sample demonstrations from the MATH Hendrycks et al. (2021) training set using 5 different random seeds, and report the average performance across them.
@@ Appendix D Results and Analysis
D.1 Training Dynamics
Figure 14 illustrates the training dynamics. For smaller models, ICPO maintains a higher policy entropy than GRPO, reflecting a broader policy support and increased exploration during training.
For larger models, their stronger ICL capability makes IEF more effective. This leads to higher policy entropy and KL divergence, while we retain the KL loss to provide a valid constraint on policy updates. As the stronger exploration, correspondingly, the generated responses become longer.
Overall, these observations indicate that ICPO facilitates stronger exploration compared to GRPO.
@@ D.1 Training Dynamics
(a) Training dynamics of Qwen3-1.7B.
(b) Training dynamics of Qwen3-8B.
Figure 14: Training dynamics of ICPO.
D.2 Pass@32 Performance
@@ D.3 OOD Performance
D.4 Generalizability Across Models
@@ E.1 Ablation on Expert Guidance
Variant MATH AIME24/25 AMC Mnrv. Avg.
GRPO 91.0 54.8 / 38.5 83.8 50.7 63.8
MATH Hendrycks et al. (2021)
ICPO 92.0 55.2 / 43.6 87.0 51.1 65.8
[…]
MathInstruct Yue et al. (2023)
ICPO 90.8 56.7 / 40.7 83.8 53.7 65.1
[…]
Table 11: Ablation analysis across different expert-domain data on Qwen3-8B.
We study the generalizability of ICPO under different sources of expert-domain data. In addition to our default setting, we adopt MathInstruct [^3] https://huggingface.co/datasets/TIGER-Lab/MathInstruct Yue et al. (2023) as an alternative expert demonstration corpus for ICPO. MathInstruct contains approximately 262k instruction–solution pairs. To enable reliable extraction of ground-truth answers for RLVR, we select a subset of 118k examples whose solutions explicitly include the phrase “The answer is”. As shown in Table 11, ICPO and its variants guided by different expert-domain data sources consistently outperform the GRPO baseline, demonstrating the robustness of our approach to the choice of expert data. Notably, although all evaluations are conducted on in-distribution benchmarks, expert data drawn from different domains exhibits varying strengths across benchmarks. This observation suggests that more carefully curated or benchmark-aligned demonstrations may further improve generalization performance across diverse tasks.

### v3

@@ Abstract
Existing Reinforcement Learning from Verifiable Rewards (RLVR) methods, such as Group Relative Policy Optimization (GRPO), have achieved remarkable progress in improving the reasoning capabilities of Large Reasoning Models (LRMs). However, they exhibit limited exploration due to reliance on on-policy rollouts which are confined to the current policy’s distribution, resulting in narrow trajectory diversity. Recent approaches attempt to expand policy coverage by incorporating trajectories generated from stronger expert models, yet this reliance increases computational cost and such advanced models are often inaccessible. To address these issues, we propose In-Context Steered Policy Optimization (ICPO), a unified framework that leverages the inherent in-context learning capability of LRMs to provide expert guidance using existing datasets. ICPO introduces mixed-policy GRPO with implicit expert forcing, which expands exploration beyond the current policy distribution without requiring advanced LRM trajectories. To further stabilize optimization, ICPO integrates expert region reject sampling to filter unreliable off-policy trajectories and annealed expert-bonus reward shaping to balance early expert guidance with later autonomous improvement. Results demonstrate that ICPO consistently enhances RLVR performance and training stability on mathematical reasoning benchmarks, revealing a scalable and effective RLVR paradigm for LRMs. Our code is available at https://github.com/Celine-hxy/ICPO.
@@
Hsiu-Yuan Huang1,2,3thanks:   Equal contribution. Work done during internship at Tencent., Chenming Tang1,211footnotemark: 1, Weijie Liu3thanks:   Corresponding author., Clive Bai3, Saiyong Yang3, Yunfang Wu1, [^2] footnotemark: 2 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University  3LLM Department, Tencent Correspondence: huang.hsiuyuan@stu.pku.edu.cn wuyf@pku.edu.cn
@@ 3 Preliminary
where $A(\cdot)$ maps demonstrations $\mathcal{D}$ to a task vector $\vartheta=A(\mathcal{D})$ that encodes the expert behavior specific to that task Hendel et al. (2023); Li et al. (2023); Todd et al. (2024); Huang et al. (2024); Liu et al. (2024); Hojel et al. (2024); Cai et al. (2025), and $\mathcal{F}(\cdot\,;\vartheta)$ represents the task-specific reasoning function that applies the $\vartheta$ to generate the prediction for query $q$. This leads to the parametric representation:
@@ 5 Experimental Setup
The complete set of hyperparameters used in our training is listed in Appendix B.1. For the main experiments, we use Qwen3-1.7B and Qwen3-8B Yang et al. (2025) as the base models and employ GRPO Shao et al. (2024) as our RL algorithm. We also include Qwen2.5-Math-7B to facilitate comparison with other related methods. We additionally evaluate ICPO on Qwen3-8B-Base and LLaMA-3.1-8B to examine its cross-model generalization. We generate a total of 8 rollout trajectories per prompt. For the on-policy baseline, we use 8 on-policy rollouts. For our mixed-policy GRPO, we follow previous work Yan et al. (2025) and use 1 off-policy rollout and 7 on-policy rollouts to ensure comparability.
Evaluation Settings.
We evaluate on six widely used mathematical reasoning benchmarks: AIME24, AIME25, AMC23 Li et al. (2024), Minerva Lewkowycz et al. (2022), Olympiad He et al. (2024), and MATH-500 Hendrycks et al. (2021). To assess generalization beyond in-domain reasoning, we further test on three out-of-distribution (OOD) benchmarks: ARC-C Clark et al. (2018), GPQA-Diamond Rein et al. (2023), and MMLU-Pro Wang et al. (2024), with multiple-choice options shuffled to prevent contamination. For main experiments, the AIME24, AIME25, and AMC23 benchmark, which have relatively small test sets, we report Avg@32 (for Pass@1), while for the other benchmarks we report Pass@1. More evaluation details are provided in Appendix B.2.
@@ 6.1 Main Results
Figure 5: Reward curves over training steps across test (Mean@2) and train sets.
Figure 6: Training dynamics of Qwen3-1.7B.
@@ 6.2 Ablation Study
Selection of Expert Data.
Method MATH AIME24/25 AMC Mnrv. Avg.
Qwen3-1.7B
GRPO 83.6 28.4 / 22.5 66.7 40.8 48.4
ICPO (PoT) 87.0 31.5 / 27.0 69.6 42.3 51.5
ICPO (CoT) 86.8 31.3 / 26.3 70.4 44.1 51.8
Qwen3-8B
GRPO 91.0 54.8 / 38.5 83.8 50.7 63.8
ICPO (PoT) 90.8 56.7 / 40.7 83.8 53.7 65.1
ICPO (CoT) 92.0 55.2 / 43.6 87.0 51.1 65.8
Table 4: Ablation on expert data selection, where CoT refers to mathematical data (MATH) and PoT refers to code-oriented OOD reasoning data.
To demonstrate the robustness of ICPO to the choice of expert data, we replace the in-domain expert data with cross-domain Program-of-Thought (PoT) data Yue et al. (2023) and conduct ICPO training under this setting. Details of data processing are provided in Appendix E.1. As shown in Table 4, ICPO consistently outperforms GRPO when using either CoT or PoT data as demonstrations for implicit expert guidance. These results suggest that ICPO is robust to the choice of expert data and highlight its potential to generalize beyond mathematical reasoning to cross-domain settings. More importantly, by changing the expert data, ICPO functions as a plug-and-play framework that flexibly reshapes the target policy distribution, enabling controllable steering of model behavior.
@@
Role of Base Model ICL Capability.
Method MATH AIME24/25 AMC Mnrv. Avg.
Pass@1
GRPO 84.6 22.5 / 21.4 71.8 43.8 48.8
$\text{GRPO}_{\text{ExtraRollouts}}$ 85.8 26.9 / 20.2 67.4 46.7 49.4
$\text{GRPO}_{\text{ExpertDomain}}$ 80.0 22.9 / 12.8 60.9 47.8 44.9
ICPO (Ours) 88.6 26.8 / 21.6 69.2 46.3 50.5
Pass@32
GRPO 96.2 53.3 / 43.3 90.0 58.8 68.3
$\text{GRPO}_{\text{ExtraRollouts}}$ 95.4 50.0 / 46.7 92.5 64.3 69.8
$\text{GRPO}_{\text{ExpertDomain}}$ 92.0 46.7 / 26.7 90.0 61.0 63.3
ICPO (Ours) 95.8 60.0 / 60.0 95.0 63.6 74.9
Table 6: Pass@1 and Pass@32 results for Qwen3-8B-Base, a model with no ICL capability.
To examine whether ICPO relies on the assumption that the base model possesses basic ICL capability, we conduct additional experiments on Qwen3-8B-Base, a model with no ICL capability. The results are summarized in Table 6. The observed Pass@1 improvement demonstrates that ICPO remains effective even when the base model lacks ICL ability. Meanwhile, the substantial Pass@32 improvement indicates that ICPO expands the exploration space rather than merely sharpening existing modes, encouraging the model to “think outside the policy.” These results show that ICPO is not predicated on the ICL capabilities of the base model Instead, ICL acts as a scaling amplifier: stronger ICL ability leads to larger improvements, yet weakly aligned models can also benefit from implicit expert forcing during RL optimization.
@@ 8 Conclusion
Our experiments focus primarily on mathematical reasoning and OOD reasoning benchmarks under the RLVR framework, including multi-domain knowledge reasoning (MMLU), open-domain QA (GPQA), and scientific reasoning tasks (ARC). Extending ICPO to other domains, such as code generation, may require task-specific adaptations, especially in reward design and evaluation. We leave a systematic investigation of these broader applications to future work.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62076008).
@@ B.2 Evaluation Settings
For main experiments, the AIME24, AIME25, and AMC23 benchmark, which have relatively small test sets, we report Avg@32 (for Pass@1), while for the other benchmarks we report Pass@1. As a complementary analysis, we also report Pass@32 results in Appendix D.1. All evaluations are conducted using the LIMO framework Ye et al. (2025). During inference, we follow Yan et al. (2025) and set the generation temperature to 0.6. For ICL evaluation, we randomly sample demonstrations from the MATH Hendrycks et al. (2021) training set using 5 different random seeds, and report the average performance across them.
@@ Appendix D Results and Analysis
D.1 Pass@32 Performance
@@ D.2 OOD Performance
D.3 Generalizability Across Models
@@ E.1 Ablation on Expert Guidance
We study the generalizability of ICPO under different sources of expert-domain data. In addition to our default setting, we adopt MathInstruct [^3] https://huggingface.co/datasets/TIGER-Lab/MathInstruct Yue et al. (2023) as an alternative expert demonstration corpus for ICPO. MathInstruct contains approximately 262k instruction–solution pairs. To enable reliable extraction of ground-truth answers for RLVR, we select a subset of 118k examples whose solutions explicitly include the phrase “The answer is”.
As shown in Table 4, ICPO consistently outperforms GRPO when using either CoT or PoT data as demonstrations for implicit expert guidance. These results suggest that ICPO is robust to the choice of expert data, and highlight its potential to generalize beyond mathematical reasoning to cross-domain settings, such as code-oriented tasks.